# Sparse factorization of the square all-ones matrix of arbitrary order

### v1

@@ 1 Introduction
The second class of solutions allows for differing factors of $J$. Among these solutions include one-peer exponential graphs (when $n=2^{q}$[13], one-peer hyper-cubes (when $n=2^{q}$[8], and $p$-peer hyper-cuboids [9, 7]. The $p$-peer hyper-cuboids serve as a factorization of $J$ for arbitrary $n$, but the sparsity of the factors depends on the prime factorization of $n$. In particular, $p$-peer hyper-cuboids are no longer sparse when the matrix order equals a large prime factor. Allowing for different factors of $J$, in general, gives greater control over the sparsity of the factors compared to the case in which all the factors are identical [7].
@@
where $J_{0}=J_{1}\oplus\cdots\oplus J_{\tau}$ with $J_{k}:=\tfrac{1}{n_{k}}\mathds{1}\mathds{1}^{\mathsf{T}}\in\mathbb{R}^{n_{k}\times n_{k}}$, $k\in[\tau]$. (Here, $\oplus$ denotes the direct sum of two matrices.) Throughout the paper, it is assumed that the partition $n=\sum_{k=1}^{\tau}n_{k}$ is given, with conditions that will be specified later (see (3)). Factorization (2) holds for arbitrary matrix order $n$ and is inspired by the applications of $J$ in decentralized averaging (and optimization). In decentralized averaging, for example, there is a group of agents where each holds a piece of information and cooperates with other agents to compute a global quantity. The communication between agents is modeled by a graph (or a sequence of graphs) ${\mathcal{G}}^{(k)}=({\mathcal{V}},W^{(k)},{\mathcal{E}}^{(k)})$. If the weight matrices $\{W^{(k)}\}$ satisfy (1), then the exact global average is computed in $q$ communication rounds. In modern application scenarios, agents can be abstracted as high-performance computing (HPC) resources and can be naturally formed into clusters [14]. Such clustering structure is captured by the proposed form of factorization (2). The block diagonal matrix $J_{0}$ models the intra-cluster communication, and each sub-block $J_{k}$, $k\in[\tau]$, can be further decomposed as (1) into, e.g., $g$-circulant matrices or $p$-peer hyper-cuboids. In contrast, the $A$-factor models the more expensive inter-cluster communication, and the main focus of this paper is to design sparse $A$-factors to reduce the communication overhead across clusters. Sparsity in $A$ is desirable in decentralized averaging (and optimization) as the communication overhead is related to the total number of nonzeros $\mathrm{nnz}(A)$ as well as the largest node degree $d_{\mathrm{max}}(A)=\max_{i}\{A_{i,:}\}$, where $A_{i,:}$ is the $i$th row of $A$.
@@
Let $\mathbb{R}$ denote the set of real numbers (i.e., scalars). Let $\mathbb{R}^{n}$ denote the set of $n$-dimensional (column) vectors. (In this paper, all vectors are column vectors.) Let $\mathbb{R}^{m\times n}$ denote the set of $m$-by-$n$ real matrices, and let $\mathbb{D}^{n}$ denote the set of $n\times n$ diagonal matrices. The set of natural numbers is denoted as :=$\mathbb{N}:=\{0,1,2,\ldots\}$, and let $\mathbb{N}_{\geq r}$ denote the set of natural numbers greater than or equal to $r\in\mathbb{N}$. For any $n\in\mathbb{N}$, let $[n]:=\{1,2,\ldots,n\}$. Let $\mathds{1}$ denote the all-ones (column) vector of compatible size. The direct sum of two matrices $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{p\times q}$ forms the block diagonal matrix $A\oplus B:=\operatorname{blkdiag}(A,B)\in\mathbb{R}^{(m+p)\times(n+q)}$.
@@ 2 Hierarchically banded matrices
$n=\sum\limits_{k=1}^{\tau}n_{k},\;\text{where}\ \{n_{k}\}_{k=1}^{\tau}\subset\mathbb{N}_{\geq 1},\ \text{and}\ n_{k}\geq\sum\limits_{j=k+1}^{\tau}n_{j}=:m_{k}\ \text{for all}\ k\in[\tau-1].$ (3)
Such a partition can be constructed systematically, e.g., via the base-$p$ representation of $n$ (with $p\in\mathbb{N}_{\geq 2}$). Overloading the binary representation, we denote the base-$p$ representation of $n$ as $(i_{\tau-1}i_{\tau-2}\cdots i_{1}i_{0})_{p}$, where $\tau=\lfloor\log_{p}(n)\rfloor+1$. Then, any integer $n\in\mathbb{N}_{\geq 2}$ can be written as $n=\sum_{k=1}^{\tau}n_{k}$, where $n_{k}=i_{\tau-k}p^{\tau-k}$. In this case, the condition $n_{k}\geq m_{k}$, for all $k\in[\tau-1]$, directly follows from the property of the base-$p$ representation. A simple example is $(n,p)=(15,2)$, and $(n_{1},n_{2},n_{3},n_{4})=(8,4,2,1)$, which follows from the binary representation $15=(1111)_{2}$.
@@
$\small\begin{bmatrix}A^{(1)}_{11}&A^{(1)}_{12}\\ \big{(}A^{(1)}_{12}\big{)}^{\mathsf{T}}&A^{(1)}_{22}\end{bmatrix}=$
@@
$\small=\begin{bmatrix}A^{(2)}_{11}&A^{(2)}_{12}\\ \big{(}A^{(2)}_{12}\big{)}^{\mathsf{T}}&A^{(2)}_{22}\end{bmatrix}$
@@
Moreover, the DSHB factorization inspires another sparse factorization (2) of $J$, which is called the sequential doubly stochastic (SDS) factorization. In this factorization, the SDS factor $A$ is doubly stochastic and additionally can be written as the product of a sequence of symmetric, doubly stochastic matrices. Although the SDS factor is not hierarchically banded (nor symmetric), it is closely related to the DSHB factorization and finds its application in decentralized optimization. The details of the SDS factorization are presented in Section 5.
@@ 3 Reduced hierarchically banded factorization
As motivated in Section 2, the (density) reduced hierarchically banded (RHB) factorization further promotes sparsity in the HB factor $A$ by requiring
$A^{(k)}_{12}[j,j]=\beta_{k}\neq 0,\ \ \text{if}\ j=1,1+n_{k+1},1+n_{k+1}+n_{k+2},\ldots,1+\sum\limits_{\ell=1}^{\tau-1}n_{k+\ell},$ (5)
i.e., only a few nonzeros exist in the diagonal entries of $A^{(k)}_{12}$, and these nonzeros are all equal to some $\beta\in\mathbb{R}$. In addition, it is also assumed that only one diagonal entry in each $A^{(k)}_{11}$ is not one:
@@ 3.1 A two-block example
where
$\alpha_{1}=\frac{n_{1}^{2}}{n}-n_{1}+1,\qquad\alpha_{2}=\frac{n_{2}^{2}}{n}-n_{2}+1,\qquad\beta=\frac{n_{1}n_{2}}{n}.$
@@ 3.2 The RHB factorization algorithm
3:Set $m_{0}\leftarrow n$.
4:for $k=1,2,\ldots,\tau-2$ do
5:     Compute the $(1,1)$-block $A^{(k)}_{11}\in\mathbb{D}^{n_{k}}$ of $A^{(k)}$:
$A^{(k)}_{11}\leftarrow\operatorname{diag}\Big{(}\tfrac{n_{k}^{2}}{m_{k-1}}-n_{k}+1,1,\ldots,1\Big{)}.$
6:     Compute the $(1,2)$-block $A^{(k)}_{12}\in\mathbb{R}^{n_{k}\times m_{k}}$:
@@
7:     Compute the $(2,2)$-block $A^{(k)}_{22}=A^{(k+1)}$ as the RHB factorization:
$\frac{1}{m_{k}}\mathds{1}_{m_{k}}\mathds{1}_{m_{k}}^{\mathsf{T}}=\mkern 1.5mu\overline{\mkern-1.5muJ}_{k}A^{(k+1)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k},$ (10)
where $\mkern 1.5mu\overline{\mkern-1.5muJ}_{k}:=J_{k+1}\oplus\cdots\oplus J_{\tau}$, and the OHHG factor $A^{(k+1)}=A^{(k)}_{22}$ is partitioned as
@@
which takes the same form as (12). We can repeat the above process for $k=1,2,\ldots,\tau-2$. When Algorithm 1 reaches iteration $k=\tau-2$, Line 7 computes the RHB factorization of the matrix
@@
$\alpha_{1}=\frac{n_{\tau-1}^{2}}{n_{\tau-1}+n_{\tau}}-n_{\tau-1}+1,\qquad\alpha_{2}=\frac{n_{\tau}^{2}}{n_{\tau-1}+n_{\tau}}-n_{\tau}+1,\qquad\beta=\frac{n_{\tau-1}n_{\tau}}{n_{\tau-1}+n_{\tau}}.$
@@
The $n\times n$ matrix $A_{\mathrm{RHB}}$ generated by Algorithm 1 is hierarchically banded, satisfies the conditions (5)–(6), and satisfies $J=J_{0}A_{\mathrm{RHB}}J_{0}$. In addition, the total number of nonzeros is $\mathrm{nnz}(A_{\mathrm{RHB}})=n+\tau(\tau-1)$, and the largest node degree is $d_{\mathrm{max}}(A_{\mathrm{RHB}})=\tau$.
@@
(The subscript in $A_{\mathrm{RHB}}$ (and $\widetilde{A}^{(k)}_{\mathrm{RHB}}$) is omitted in the proof for readability.) The hierarchically banded structure of $A$ follows from the recursive nature of Algorithm 1, and in particular, the recursive partition of $A^{(k)}$ in 7. Similarly, $A$ satisfies the conditions (5)–(6) due to the assignment of values in $A^{(k)}_{11}$ and $A^{(k)}_{12}$ in 5 and 6. Next, the factorization $J=J_{0}AJ_{0}=J_{0}A^{(1)}J_{0}$ holds by recursively applying (10) for $k=\tau-1,\tau-2,\ldots,1$. Finally, one has for all $k\in[\tau-1]$ that $\mathrm{nnz}(A^{(k)}_{11})=n_{k}$, $\mathrm{nnz}(A^{(k)}_{12})=\tau-k$, and $\mathrm{nnz}(A^{(\tau)})=n_{\tau}$. So, the total number of nonzeros is
@@ 4.1 A two-block example
$J_{1}A_{12}J_{2}=\frac{1}{n_{1}n_{2}}\mathds{1}_{n_{1}}\Big{(}\mathds{1}_{n_{1}}^{\mathsf{T}}A_{12}\mathds{1}_{n_{2}}\Big{)}\mathds{1}_{n_{2}}^{\mathsf{T}}=\frac{\beta}{n_{1}}.$
@@
With the subblock $A_{12}$ settled, the doubly stochastic property of $A$ implies that the subblocks $A_{11}\in\mathbb{D}^{n_{1}}$ and $A_{22}\in\mathbb{D}^{n_{2}}$ are diagonal matrices satisfying
@@
$A=\left[\begin{array}[]{cc|c}\tfrac{n_{2}}{n}I_{n_{2}}&0&\tfrac{n_{1}}{n}I_{n_{2}}\\ 0&I_{n_{1}-n_{2}}&0\\ \hline\cr\tfrac{n_{1}}{n}I_{n_{2}}&0&\tfrac{n_{2}}{n}I_{n_{1}}\end{array}\right].$ (15)
@@ 4.2 The DSHB factorization algorithm
The $n\times n$ matrix $A_{\mathrm{DSHB}}$ generated by Algorithm 2 is doubly stochastic, hierarchically banded, and satisfies $J=J_{0}A_{\mathrm{DSHB}}J_{0}$. Each matrix in the scaled HB sequence $\{\widetilde{A}^{(k)}_{\mathrm{DSHB}}\}_{k=1}^{\tau}$ generated by Algorithm 2 is doubly stochastic. In addition, $\mathrm{nnz}(A_{\mathrm{DSHB}})=\sum_{k=1}^{\tau}kn_{k}$, and $d_{\mathrm{max}}(A_{\mathrm{DSHB}})=\tau$.
@@
(The subscript in $A_{\mathrm{DSHB}}$ (and $\widetilde{A}^{(k)}_{\mathrm{DSHB}}$) is omitted in the proof for readability.) The doubly stochastic property of $A$ and $\{\widetilde{A}^{(k)}\}$ follows from the assignments (16)–(17) and condition (18). The hierarchically banded structure of $A$ follows from the recursive nature of Algorithm 2, and in particular, the recursive partition of $\widetilde{A}^{(k)}$ in 7. Next, the factorization $J=J_{0}AJ_{0}$ holds by recursively applying (18) for $k=\tau-1,\tau-2,\ldots,1$. Finally, the number of nonzeros and the largest node degree can be calculated using the same approach as for the RHB factor in Theorem 1. ∎

### v2

@@ 1 Introduction
The second class of solutions allows for differing factors of $J$. Among these solutions include one-peer exponential graphs (when $n=2^{q}$[21], one-peer hyper-cubes (when $n=2^{q}$[14], $p$-peer hyper-cuboids [17, 12], and deformable butterfly (DeBut) matrices [10]. Remarkably, the butterfly matrices [4], which were originally proposed for more general linear transforms and used in deep neural networks, reduce to one-peer hyper-cubes when we study the factorization (1) with $n=2^{q}$. As extensions of the one-peer hyper-cubes (and butterfly matrices), the $p$-peer hyper-cuboids and the DeBut matrices serve as factorization of $J$ for arbitrary $n$, and the sparsity of both factors depends on the prime factorization of $n$. In particular, when $n$ is a large primal number, both $p$-peer hyper-cuboids and the DeBut matrices reduce to the fully dense matrix $J$. Allowing for different factors of $J$, in general, gives greater control over the sparsity of the factors compared to the case in which all the factors are identical [12].
@@
where $J_{0}=J_{1}\oplus\cdots\oplus J_{\tau}$ with $J_{k}:=\tfrac{1}{n_{k}}\mathds{1}\mathds{1}^{\mathsf{T}}\in\mathbb{R}^{n_{k}\times n_{k}}$, $k\in[\tau]$. (Here, $\oplus$ denotes the direct sum of two matrices.) Throughout the paper, it is assumed that the partition $n=\sum_{k=1}^{\tau}n_{k}$ is given, with conditions that will be specified later (see (3)). Factorization (2) holds for arbitrary matrix order $n$ and is inspired by the applications of $J$ in decentralized averaging (and optimization). In decentralized averaging, for example, a group of agents each holds a piece of information and cooperates with other agents to compute a global quantity. The communication between agents is modeled by a graph (or a sequence of graphs) $\mathcal{G}^{(k)}=(\mathcal{V},W^{(k)},\mathcal{E}^{(k)})$. If the weight matrices $\{W^{(k)}\}$ satisfy (1), then the exact global average is computed in $q$ communication rounds. In modern application scenarios, agents can be abstracted as high-performance computing (HPC) resources and naturally formed into clusters [9, 1, 22]; see Section 6 for a more detailed discussion. Such clustering structure is captured by the proposed form of factorization (2). The block diagonal matrix $J_{0}$ models the intra-cluster communication, and each sub-block $J_{k}$, $k\in[\tau]$, can be further decomposed as (1) into, e.g., $p$-peer hyper-cuboids. In contrast, the $A$-factor models the more expensive inter-cluster communication, and the main focus of this paper is to design sparse $A$-factors to reduce the communication overhead across clusters. Sparsity in $A$ is desirable in decentralized averaging (and optimization) as the communication overhead is related to the total number of nonzeros $\mathrm{nnz}(A)$ as well as the largest node degree $d_{\mathrm{max}}(A)=\max_{i}\{\mathrm{nnz}(A_{i,:})\}$, where $A_{i,:}$ is the $i$th row of $A$.
@@
[…]
@@ 2 Hierarchically banded matrices
$n=\sum_{k=1}^{\tau}n_{k},\;\text{where}\ \{n_{k}\}_{k=1}^{\tau}\subset\mathbb{N}_{\geq 1},\ \text{and}\ n_{k}\geq\sum_{j=k+1}^{\tau}n_{j}=:m_{k}\ \text{for all}\ k\in[\tau-1].$ (3)
Such a partition can be constructed systematically, e.g., via the base-$p$ representation of $n$ (with $p\in\mathbb{N}_{\geq 2}$). Overloading the binary representation, we denote the base-$p$ representation of $n$ as $(i_{\tau-1}i_{\tau-2}\cdots i_{1}i_{0})_{p}$, where $\tau=\lfloor\log_{p}(n)\rfloor+1$. Then, any integer $n\in\mathbb{N}_{\geq 2}$ can be written as $n=\sum_{k=1}^{\tau}n_{k}$, where $n_{k}=i_{\tau-k}p^{\tau-k}$. (For those $i_{k}$’s being zero, they are removed before the construction of the partition.) By construction, the condition $n_{k}\geq m_{k}$, for all $k\in[\tau-1]$, directly follows from the property of the base-$p$ representation. A simple example is $(n,p)=(15,2)$, and $(n_{1},n_{2},n_{3},n_{4})=(8,4,2,1)$, which follows from the binary representation $15=(1111)_{2}$.
@@
$\begin{bmatrix}A^{(1)}_{11}&A^{(1)}_{12}\\ \big{(}A^{(1)}_{12}\big{)}^{\mathsf{T}}&A^{(1)}_{22}\end{bmatrix}=\;$
@@
$\;=\begin{bmatrix}A^{(2)}_{11}&A^{(2)}_{12}\\ \big{(}A^{(2)}_{12}\big{)}^{\mathsf{T}}&A^{(2)}_{22}\end{bmatrix}$
@@
Moreover, the DSHB factorization inspires another sparse factorization (2) of $J$, which is called the sequential doubly stochastic (SDS) factorization. In this factorization, the SDS factor $A$ is doubly stochastic and can be written as the product of a sequence of symmetric, doubly stochastic matrices. Although the SDS factor is not hierarchically banded (nor symmetric), it is closely related to the DSHB factorization and finds its application in decentralized optimization. The details of the SDS factorization are presented in Section 5.
@@ 3 Reduced hierarchically banded factorization
As discussed in Section 2, the (density) reduced hierarchically banded (RHB) factorization further promotes sparsity in the HB factor $A$ by requiring
$A^{(k)}_{12}[j,j]\neq 0,\ \ \text{if}\ j=1,\ 1+n_{k+1},\ 1+n_{k+1}+n_{k+2},\ \ldots,\ 1+\sum_{\ell=1}^{\tau-k-1}n_{k+\ell},$ (5)
i.e., only a few nonzeros exist in the diagonal entries of $A^{(k)}_{12}$. In addition, it is also assumed that only one diagonal entry in each $A^{(k)}_{11}$ is not one:
@@ 3.1 A two-block example
where $\alpha_{1}=\frac{n_{1}^{2}}{n}-n_{1}+1$, $\alpha_{2}=\frac{n_{2}^{2}}{n}-n_{2}+1$, and $\beta=\frac{n_{1}n_{2}}{n}$.
@@ 3.2 The RHB factorization algorithm
3:for $k=1,2,\ldots,\tau-2$ do
4:     Compute the $(1,1)$-block $A^{(k)}_{11}\in\mathbb{D}^{n_{k}}$ of $A^{(k)}$:
$A^{(k)}_{11}\leftarrow\operatorname{diag}\Big{(}\tfrac{n_{k}^{2}}{n}-n_{k}+1,1,\ldots,1\Big{)}.$
5:     Compute the $(1,2)$-block $A^{(k)}_{12}\in\mathbb{R}^{n_{k}\times m_{k}}$:
@@
6:     Compute the $(2,2)$-block $A^{(k)}_{22}=A^{(k+1)}$ as the RHB factorization:
$\tfrac{1}{m_{k}}\mathds{1}_{m_{k}}\mathds{1}_{m_{k}}^{\mathsf{T}}=\mkern 1.5mu\overline{\mkern-1.5muJ}_{k}A^{(k+1)}\mkern 1.5mu\overline{\mkern-1.5muJ}_{k},$ (10)
where $\mkern 1.5mu\overline{\mkern-1.5muJ}_{k}:=J_{k+1}\oplus\cdots\oplus J_{\tau}$, and the RHB factor $A^{(k+1)}=A^{(k)}_{22}$ is partitioned as
@@
which takes the same form as (12). We can then repeat the above process for $k=1,2,\ldots,\tau-2$. When Algorithm 1 reaches iteration $k=\tau-2$, Line 6 computes the RHB factorization of the matrix
@@
$\alpha_{1}=\frac{n_{\tau-1}^{2}}{n}-n_{\tau-1}+1,\qquad\alpha_{2}=\frac{n_{\tau}^{2}}{n}-n_{\tau}+1,\qquad\beta=\frac{n_{\tau-1}n_{\tau}}{n}.$
@@
The $n\times n$ matrix $A_{\mathrm{RHB}}$ generated by Algorithm 1 is hierarchically banded and satisfies $J=J_{0}A_{\mathrm{RHB}}J_{0}$ as well as conditions (5)–(6). In addition, the total number of nonzeros is $\mathrm{nnz}(A_{\mathrm{RHB}})=n+\tau(\tau-1)$, and the largest node degree is $d_{\mathrm{max}}(A_{\mathrm{RHB}})=\tau$.
@@
(The subscript in $A_{\mathrm{RHB}}$ is omitted in the proof for readability.) The hierarchically banded structure of $A$ follows from the recursive nature of Algorithm 1, and in particular, the recursive partition of $A^{(k)}$ in Line 6 of Algorithm 1. Similarly, $A$ satisfies the conditions (5)–(6) due to the assignment of values in $A^{(k)}_{11}$ and $A^{(k)}_{12}$ in Lines 4 and 5. Next, the factorization $J=J_{0}AJ_{0}=J_{0}A^{(1)}J_{0}$ holds by recursively applying (10) for $k=\tau-1,\tau-2,\ldots,1$. Finally, one has for all $k\in[\tau-1]$ that $\mathrm{nnz}(A^{(k)}_{11})=n_{k}$, $\mathrm{nnz}(A^{(k)}_{12})=\tau-k$, and $\mathrm{nnz}(A^{(\tau)})=n_{\tau}$. So, the total number of nonzeros is
@@ 4.1 A two-block example
$J_{1}A_{12}J_{2}=\frac{1}{n_{1}n_{2}}\mathds{1}_{n_{1}}\Big{(}\mathds{1}_{n_{1}}^{\mathsf{T}}A_{12}\mathds{1}_{n_{2}}\Big{)}\mathds{1}_{n_{2}}^{\mathsf{T}}=\frac{\beta}{n_{1}}\mathds{1}_{n_{1}}\mathds{1}_{n_{2}}^{\mathsf{T}}.$
@@
With the sub-block $A_{12}$ settled, the doubly stochastic property of $A$ implies that the sub-blocks $A_{11}\in\mathbb{D}^{n_{1}}$ and $A_{22}\in\mathbb{D}^{n_{2}}$ are diagonal matrices satisfying
@@
$A=\left[\begin{array}[]{cc|c}\tfrac{n_{2}}{n}I_{n_{2}}&0&\tfrac{n_{1}}{n}I_{n_{2}}\\ 0&I_{n_{1}-n_{2}}&0\\ \hline\cr\tfrac{n_{1}}{n}I_{n_{2}}&0&\tfrac{n_{2}}{n}I_{n_{2}}\end{array}\right].$ (15)
@@ 4.2 The DSHB factorization algorithm
The $n\times n$ matrix $A_{\mathrm{DSHB}}$ generated by Algorithm 2 is doubly stochastic, hierarchically banded, and satisfies $J=J_{0}A_{\mathrm{DSHB}}J_{0}$. Each matrix in the scaled HB sequence $\{\widetilde{A}^{(k)}_{\mathrm{DSHB}}\}_{k=1}^{\tau}$ generated by Algorithm 2 is doubly stochastic. In addition, it holds that
$\mathrm{nnz}(A_{\mathrm{DSHB}})=\sum_{k=1}^{\tau}kn_{k},\qquad d_{\mathrm{max}}(A_{\mathrm{DSHB}})=\tau.$
@@
(The subscript in $A_{\mathrm{DSHB}}$ (and $\widetilde{A}^{(k)}_{\mathrm{DSHB}}$) is omitted in the proof for readability.) The doubly stochastic property of $A$ and $\{\widetilde{A}^{(k)}\}$ follows from the assignments (16)–(17) and condition (18). The hierarchically banded structure of $A$ follows from the recursive nature of Algorithm 2, and in particular, the recursive partition of $\widetilde{A}^{(k)}$ in Line 7. Next, the factorization $J=J_{0}AJ_{0}$ holds by recursively applying (18) for $k=\tau-1,\tau-2,\ldots,1$. Finally, the number of nonzeros and the largest node degree can be calculated using the same approach as for the RHB factor in Theorem 1. ∎